Grade 11Algebra

A grocer buys two kinds of wheat A and B. The rate of A is rupees x per kg and rate of B is rupees y per kg .He mixes both wheat and obtained the wheat with rate Rupees z per kg then ratio of the two varieties A and B of wheat in the mixture is
1.(y+z):(z+x)
2.(z-x):(y-z)
3.x:y
4.(y-z):(z-x)

Rankush
8 Years agoGrade 11

1 Answer

Mujtaba Basheer
8 Years ago
Let the quantity of wheat A purchased be a kg and the quantity of wheat B purchased be b kg.
Hence the rate of the mixed wheat is (ax+by)/(a+b) which, according to the question, is equal to z.
Hence, (ax+by)/(a+b)=z
             ax+by=az+bz
             b(y-z)=a(z-x)
Hence, a:b=(y-z):(z-x)
So the fourth option is correct.